OpenStudy (vshiroky):

at the beginning of the bake sale, mary set aside 4 pies for herself. karen then bought half of the remaining pies for her birthday party, and 5 more pies were sold during the sale. when the sale was over there was 16 pies remaining. how many pies were there before the bake sale started?

OpenStudy (vshiroky):

@johnweldon1993 can you help me please :) it has been a long time since I have done this

OpenStudy (vshiroky):

I'll give you a medal for helping someone please :)

OpenStudy (vshiroky):

@jdoe0001

OpenStudy (anonymous):

I believe it should be 50

OpenStudy (jdoe0001):

the original amount they had was "x" we know they sold 4 to Mary, then after they sold that to Mary, they're left with " x -4 " then Karen came in and bought half that, that is \(\bf \cfrac{x-4}{2}\) and after that, 5 more pies were sold so let's see how many they sold \(\bf (x-4)+\cfrac{x-4}{2}+5\) they had 16 pies leftover so we could say that the total amount of pies is \(\bf (x-4)+\cfrac{x-4}{2}+5+{\color{red}{ 16}}\) and that should equals "x" solve for "x"

OpenStudy (anonymous):

sorry 46, add 5 to 16=21 double that 42 add the pies that were set aside + 4 = 46

OpenStudy (jdoe0001):

hmme lemme recheck myself

OpenStudy (anonymous):

is 46 one of the answers? @vshiroky

OpenStudy (vshiroky):

It doesn't give options

OpenStudy (jdoe0001):

I think it should be instead just 4 for Mary

OpenStudy (vshiroky):

It was 46. Thank you :)

OpenStudy (jdoe0001):

\(\bf 4+\cfrac{x-4}{2}+5+{\color{red}{ 16}}=x\) keep in mind that when Mary bought 4 from the total "x", she left only x-4 and Karen got half that, that is \(\bf \cfrac{x-4}{2}\)

OpenStudy (anonymous):

it should be 46, because there were 16 pies remaining, someone bought 5 pies, 5+16=21 someone bought half the remaining pies after the 4 were set aside 21x2=42 4 pies were set aside 42+4=46

OpenStudy (anonymous):

ya np @vshiroky